# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ Appendix C Development tools, evaluation approaches, and evaluation metrics
# System Semantic Web languages, standards, reasoners, and development frameworks Other programming languages, libraries, and frameworks Database systems, data repositories, and data analysis tools Evaluation approaches Evaluation metrics
1 Akhtar et al. [76] Protégé NetLogo None mentioned Case study; System simulation None mentioned
@@
4 Ali et al. [79] Apache Jena; OWL; Pellet; Protégé; SPARQL; SWRL Java None mentioned Comparison with SOTA; Expert evaluation of system; Ontology validation Accuracy; Precision; Recall; F-score
5 Alti et al. [80] Apache Jena; Protégé; SWRL JADE; Java; Java Expert System Shell MySQL Case study; Comparison with SOTA Execution time; Optimality; Application lifetime; No. of discovered services
6 Chatterjee et al. [81] Apache Jena; Apache Jena Fuseki; HermiT; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Ontology validation Ontology reasoning time
@@
8 De Brouwer et al. [83] Apache Jena; Protégé; RDF; SPARQL None mentioned MongoDB System queries; Use case scenarios None mentioned
9 El-Sappagh et al. [84] HermiT; OWL; Pellet; Protégé; SPARQL; SWRL SQL SQLite Ontology validation; Comparison with SOTA; Comparison with expert opinion; Use case scenarios Correctness; Completeness; Extensibility; Conciseness; Organisational fitness
@@
11 Esposito et al. [86] OWL; Pellet; Protégé Java SQLite Case study; ALMA method; Ontology validation Modifiability
12 Fenza et al. [87] OWL; SPARQL JADE None mentioned Case study; Comparison with logic-based matching Precision; Recall
@@
14 Hadjadj and Halimi [89] Apache Jena; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Use case scenario; Comparison with expert opinion Similarity between system and expert opinion
@@
17 Hristoskova et al. [92] OWL; Pellet; SWRL None mentioned None mentioned Expert evaluation; User evaluation; Ontology validation Performance; Scalability
18 Hussain and Park [93] Protégé; RDF None mentioned Apache ActiveMQ; Apache Hadoop; Apache Spark; Elasticsearch; MariaDB Comparison with SOTA AUC; Accuracy; Precision; Recall; Neg. predictive value
[…]
Table continued on next page.
Summary of development tools, evaluation approaches, and evaluation metrics used by the systems.
# System Semantic Web languages, standards, reasoners, and development frameworks Other programming languages, libraries, and frameworks Database systems, data repositories, and data analysis tools Evaluation approaches Evaluation metrics
[…]
@@
22 Kim et al. [97] Apache Jena; OWL; Protégé SQL None mentioned Ontology validation; User evaluation Precision; Recall; F-score; Likert score
@@
24 Lopes de Souza et al. [99] RDF GraphQL; Java; Javascript; Node.js; React SQLite Ontology evaluation None mentioned
25 Mavropoulos et al. [100] OWL; SPARQL None mentioned GraphDB Comparison with SOTA; End user evaluation Accuracy; Precision; Recall; F-score; Likert scale
26 Mcheick et al. [101] None mentioned Netica None mentioned Use case scenarios Adaptability
27 Mezghani et al. [102] Apache Jena Fuseki; RDF; SPARQL SQL None mentioned Use case scenario None mentioned
28 Minutolo et al. [103] OWL; SPARQL None mentioned None mentioned Case study None mentioned
@@
33 Stavropoulos et al. [74] OWL; SHACL; SPARQL Python; Django GraphDB Use case scenarios; Focus group with clinicians Scalability; Likert scale
@@
36 Villarreal et al. [109] None mentioned Java MySQL Case study; ALMA method; User evaluation Response time; Usability; Recommendation suitability
@@
38 Yu et al. [70] None mentioned Python; spaCy; NetworkX None mentioned ML model performance; Use case scenarios AUC; Chronic Care Model criteria
@@
40 Zafeiropoulos et al. [75] OWL; Pellet; Protégé; RDF; SPARQL; SWRL Owlready2; Python; PyTorch; RDFlib None mentioned Use case scenario; Ontology evelaution; ML model performance MAE; MSE; R-Squared
@@
ALMA - Architecture-Level Modifiability Analysis; AUC - Area Under the Curve; BN - Bayesian Network; CCM - Chronic Care Model; JADE - Java Agent DEvelopment Framework; MAE - Mean Absolute Error; ML - Machine Learning; OWL - Web Ontology Language; RDF - Resource Description Framework; RML - RDF Mapping Language; (R)MSE - (Root) Mean Square Error; SHACL - Shapes Constraint Language; SOTA - State of the Art; SPARQL - SPARQL Protocol and RDF Query Language; SWRL - Semantic Web Rule Language; WEKA - Waikato Environment for Knowledge Analysis
@@ Appendix D Quality assessment
# System Year Data & devices (score out of 3) System & components development (score out of 3) Rigour of evaluation (score out of 7) Accessibility of research outputs (score out of 2) Overall (score out of 15)
1 Akhtar et al. [76] 2022 ✗ (0) Medium (2) Low (1) ✗ (0) Low (3)
2 Ali et al. [77] 2021 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
3 Ali et al. [78] 2020 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
4 Ali et al. [79] 2018 Low (1) Medium (2) High (6) ✗ (0) Medium (9)
5 Alti et al. [80] 2022 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
6 Chatterjee et al. [81] 2021 Medium (2) Very High (3) Medium (3) Very High (2) Medium (10)
7 Chiang and Liang [82] 2015 Low (1) Very High (3) Low (1) ✗ (0) Low (5)
8 De Brouwer et al. [83] 2022 Very High (3) Medium (2) Low (2) Medium (1) Medium (8)
9 El-Sappagh et al. [84] 2019 ✗ (0) Very High (3) Medium (4) Medium (1) Medium (8)
10 Elhadj et al. [85] 2021 ✗ (0) Very High (3) Medium (3) ✗ (0) Low (6)
11 Esposito et al. [86] 2018 Very High (3) Medium (2) High (5) ✗ (0) Medium (10)
12 Fenza et al. [87] 2012 ✗ (0) Medium (2) Low (2) ✗ (0) Low (4)
13 Garcia-Valverde et al. [88] 2014 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
14 Hadjadj and Halimi [89] 2021 Medium (2) Very High (3) Medium (3) ✗ (0) Medium (8)
15 Henaien et al. [90] 2020 Low (1) Medium (2) ✗ (0) ✗ (0) Low (3)
16 Hooda and Rani [91] 2020 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
17 Hristoskova et al. [92] 2014 Very High (3) Medium (2) Medium (3) ✗ (0) Medium (8)
18 Hussain and Park [93] 2021 Very High (3) Medium (2) Medium (4) ✗ (0) Medium (9)
[…]
21 Khozouie et al. [96] 2018 ✗ (0) Medium (2) Low (2) ✗ (0) Low (4)
22 Kim et al. [97] 2014 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
23 Kordestani et al. [98] 2021 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
24 Lopes de Souza et al. [99] 2023 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
25 Mavropoulos et al. [100] 2021 Very High (3) Low (1) High (6) ✗ (0) Medium (10)
26 Mcheick et al. [101] 2016 Low (1) Low (1) Low (2) ✗ (0) Low (4)
27 Mezghani et al. [102] 2015 ✗ (0) Low (1) Low (2) ✗ (0) Low (3)
28 Minutolo et al. [103] 2016 ✗ (0) Medium (2) Medium (3) ✗ (0) Low (5)
29 Peral et al. [104] 2018 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
30 Reda et al. [73] 2022 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
31 Rhayem et al. [105] 2021 Medium (2) Medium (2) High (5) ✗ (0) Medium (9)
32 Spoladore et al. [106] 2021 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
33 Stavropoulos et al. [74] 2021 Very High (3) Medium (2) Medium (3) ✗ (0) Medium (8)
34 Titi et al. [107] 2019 ✗ (0) Very High (3) Low (1) ✗ (0) Low (4)
35 Vadillo et al. [108] 2013 Low (1) Very High (3) Medium (3) ✗ (0) Medium (7)
36 Villarreal et al. [109] 2014 Medium (2) Low (1) High (6) ✗ (0) Medium (9)
37 Xu et al. [72] 2017 ✗ (0) Low (1) Medium (3) ✗ (0) Low (4)
38 Yu et al. [70] 2022 Medium (2) Medium (1) High (5) ✗ (0) Medium (8)
39 Yu et al. [110] 2017 Low (1) Very High (3) Low (1) ✗ (0) Low (5)
40 Zafeiropoulos et al. [75] 2024 Medium (2) Very High (3) Medium (3) Very High (2) Medium (10)
41 Zeshan et al. [111] 2023 Low (1) Medium (2) Medium (3) ✗ (0) Low (6)
42 Zhang et al. [112] 2014 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
43 Zhou et al. [71] 2022 ✗ (0) ✗ (0) Low (1) ✗ (0) Low (1)

### v3

@@ Appendix C Development tools, evaluation approaches, and evaluation metrics
# System Semantic Web-based languages, standards, reasoners, and development frameworks Other programming languages, libraries, and frameworks Database systems, data repositories, and data analysis tools Evaluation approaches Evaluation metrics
1 Akhtar et al. [109] Protégé; SWRL NetLogo None mentioned Case study; System simulation None mentioned
@@
4 Ali et al. [112] Apache Jena; OWL; Pellet; Protégé; SPARQL; SWRL Java None mentioned User evaluation with real users; Comparison with SOTA; Expert evaluation of system; Ontology validation; System queries Accuracy; Precision; Recall; F-score
5 Alti et al. [113] Apache Jena; Protégé; SWRL JADE; Java; Java Expert System Shell MySQL Case study; Comparison with SOTA; NFR validation Execution time; Optimality; Application lifetime; No. of discovered services
6 Ammar et al. [108] LDflex; OWL; RDF; RDFlib; Solid; SPARQL JavaScript; Python None mentioned Use case scenarios; Comparison with SOTA Utilization of public health library; incorporation of observations of daily living; incorporation of social determinants of health
7 Bampi et al. [87] Apache Jena; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Use case scenario; System simulation; NFR validation System latency; resource usage (memory & CPU); Comparison with SOTA
8 Chatterjee et al. [114] Apache Jena; Apache Jena Fuseki; HermiT; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Ontology validation; System queries Ontology reasoning time
@@
10 De Brouwer et al. [115] Apache Jena; Protégé; RDF; SPARQL None mentioned MongoDB System queries; Use case scenarios with real users None mentioned
11 El-Sappagh et al. [116] HermiT; OWL; Pellet; Protégé; SPARQL; SWRL SQL SQLite Ontology validation; Comparison with SOTA; Comparison with expert opinion; Use case scenarios; NFR validation; System queries Correctness; Completeness; Extensibility; Conciseness; Organisational fitness
@@
13 Esposito et al. [90] OWL; Pellet; Protégé; SWRL Java SQLite Case study with real users; ALMA method; Ontology validation; NFR validation Modifiability
14 Fenza et al. [117] OWL; SPARQL JADE; MATLAB; Fuzzy Control Language None mentioned Case study; Comparison with logic-based matching Precision; Recall
15 Garcia-Moreno et al. [91] None mentioned None mentioned None mentioned ML model performance; NFR validation; User evaluation with real users Accuracy; F-score; Sensitivity; Specificity; Reusability; Extensibility; Adaptability
@@
17 Hadjadj and Halimi [118] Apache Jena; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Use case scenario; Comparison with expert opinion; System queries Similarity between system and expert opinion
@@
20 Hristoskova et al. [119] OWL; Pellet; SWRL None mentioned None mentioned Expert evaluation; User evaluation with real users; Ontology validation; NFR validation Performance; Scalability
21 Hussain and Park [120] Protégé; RDF None mentioned Apache ActiveMQ; Apache Hadoop; Apache Spark; Elasticsearch; MariaDB Comparison with SOTA; User evaluation with real users AUC; Accuracy; Precision; Recall; Neg. predictive value
22 Ivașcu and Negru [95] Apache Jena; Apache Jena Fuseki; Protégé; SPARQL JADE; WEKA None mentioned Comparison with SOTA; System simulation Accuracy; Precision; Recall; F-score
23 Ivașcu et al. [96] Apache Jena; Protégé None mentioned None mentioned Use case scenarios None mentioned
@@
25 Kilintzis et al. [97] Apache Jena Fuseki; OWL; RDF; SPARQL; SPIN None mentioned Virtuoso User evaluation with real users; NFR validation; System queries Robustness; Data integrity
26 Kim et al. [122] Apache Jena; OWL; Protégé SQL None mentioned Ontology validation; User evaluation; System queries with real users Precision; Recall; F-score; Likert scale
@@
28 Lopes de Souza et al. [124] RDF; SWRL; SPARQL GraphQL; Java; Javascript; Node.js; React SQLite Ontology evaluation None mentioned
29 Martella et al. [125] Protégé; SPARQL Java; Python Apache Spark; MongoDB; PostgreSQL Case study with simulated users; NFR validation; System simulation Response time; Throughput; Delay threshold
30 Mavropoulos et al. [126] OWL; SPARQL None mentioned GraphDB Comparison with SOTA; User evaluation with real users Accuracy; Precision; Recall; F-score; Likert scale
31 Mcheick et al. [127] None mentioned Netica None mentioned Use case scenarios; NFR validation Adaptability
32 Mezghani et al. [128] Apache Jena Fuseki; RDF; SPARQL; SWRL SQL None mentioned Use case scenario None mentioned
33 Minutolo et al. [129] OWL; SPARQL; SWRL None mentioned None mentioned Case study None mentioned
@@
38 Stavropoulos et al. [99] OWL; SHACL; SPARQL Python; Django GraphDB Use case scenarioswith real users; NFR validation; Focus group with clinicians Scalability; Likert scale
@@
41 Villarreal et al. [102] None mentioned Java MySQL Case study; ALMA method; User evaluation with real users; NFR validation Response time; Usability; Recommendation suitability
@@
43 Yu et al. [104] None mentioned Python; spaCy; NetworkX None mentioned ML model performance; Use case scenarios with real users AUC; Chronic Care Model criteria
@@
45 Zafeiropoulos et al. [107] OWL; Pellet; Protégé; RDF; RDFlib; SPARQL; SWRL; Owlready2 Python; PyTorch None mentioned Use case scenario; Ontology evaluation; System queries; ML model performance MAE; MSE; R-Squared
@@
ALMA - Architecture-Level Modifiability Analysis; AUC - Area Under the Curve; BN - Bayesian Network; CCM - Chronic Care Model; JADE - Java Agent DEvelopment Framework; MAE - Mean Absolute Error; ML - Machine Learning; NFR - non-functional requirements; OWL - Web Ontology Language; RDF - Resource Description Framework; RML - RDF Mapping Language; (R)MSE - (Root) Mean Square Error; SHACL - Shapes Constraint Language; Solid - Social Linked Data; SOTA - State of the Art; SPARQL - SPARQL Protocol and RDF Query Language; SWRL - Semantic Web Rule Language; WEKA - Waikato Environment for Knowledge Analysis
@@ Appendix D Quality assessment
# System Year Data & devices (score out of 3) System & components development (score out of 3) Rigour of evaluation (score out of 7) Accessibility of research outputs (score out of 2)
1 Akhtar et al. [109] 2022 ✗ (0) Medium (2) Low (1) ✗ (0)
2 Ali et al. [110] 2021 Medium (2) Medium (2) Medium (4) ✗ (0)
3 Ali et al. [111] 2020 Medium (2) Medium (2) Medium (4) ✗ (0)
4 Ali et al. [112] 2018 Low (1) Medium (2) High (6) ✗ (0)
5 Alti et al. [113] 2021 Low (1) Medium (2) Low (2) ✗ (0)
6 Ammar et al. [108] 2021 Low (1) Medium (2) Medium (3) ✗ (0)
7 Bampi et al. [87] 2025 High (3) Medium (2) Low (2) High (2)
8 Chatterjee et al. [114] 2021 Medium (2) High (3) Medium (3) High (2)
9 Chiang and Liang [88] 2015 Low (1) High (3) Low (1) ✗ (0)
10 De Brouwer et al. [115] 2022 High (3) Medium (2) Low (2) Medium (1)
11 El-Sappagh et al. [116] 2019 ✗ (0) High (3) Medium (4) Medium (1)
12 Elhadj et al. [89] 2021 ✗ (0) High (3) Medium (3) ✗ (0)
13 Esposito et al. [90] 2018 High (3) Medium (2) High (5) ✗ (0)
14 Fenza et al. [117] 2012 ✗ (0) Medium (2) Low (2) ✗ (0)
15 Garcia-Moreno et al. [91] 2023 High (3) Low (1) Medium (4) ✗ (0)
16 Garcia-Valverde et al. [92] 2014 Medium (2) Low (1) Low (2) ✗ (0)
17 Hadjadj and Halimi [118] 2021 Medium (2) High (3) Medium (3) ✗ (0)
18 Henaien et al. [93] 2020 Low (1) Medium (2) ✗ (0) ✗ (0)
19 Hooda and Rani [94] 2020 Low (1) Medium (2) Low (1) ✗ (0)
20 Hristoskova et al. [119] 2014 High (3) Medium (2) Medium (3) ✗ (0)
21 Hussain and Park [120] 2021 High (3) Medium (2) Medium (4) ✗ (0)
22 Ivașcu and Negru [95] 2021 High (3) High (3) Medium (3) ✗ (0)
23 Ivașcu et al. [96] 2015 ✗ (0) Medium (2) ✗ (0) ✗ (0)
24 Khozouie et al. [121] 2018 ✗ (0) Medium (2) Low (2) ✗ (0)
25 Kilintzis et al. [97] 2019 Medium (2) High (3) Medium (4) Medium (1)
26 Kim et al. [122] 2014 Medium (2) Medium (2) Medium (4) ✗ (0)
27 Kordestani et al. [123] 2021 Medium (2) Low (1) Low (2) ✗ (0)
28 Lopes de Souza et al. [124] 2023 Low (1) Medium (2) Low (1) ✗ (0)
29 Martella et al. [125] 2025 High (3) High (3) Medium (3) ✗ (0)
30 Mavropoulos et al. [126] 2021 High (3) Low (1) High (6) ✗ (0)
31 Mcheick et al. [127] 2016 Low (1) Low (1) Low (2) ✗ (0)
32 Mezghani et al. [128] 2015 ✗ (0) Low (1) Low (2) ✗ (0)
33 Minutolo et al. [129] 2016 ✗ (0) Medium (2) Medium (3) ✗ (0)
34 Peral et al. [130] 2018 Medium (2) Low (1) Low (2) ✗ (0)
35 Reda et al. [106] 2022 Low (1) Medium (2) Low (1) ✗ (0)
36 Rhayem et al. [83] 2021 Medium (2) Medium (2) High (5) ✗ (0)
37 Spoladore et al. [98] 2021 Low (1) Medium (2) Low (2) ✗ (0)
38 Stavropoulos et al. [99] 2021 High (3) Medium (2) Medium (3) ✗ (0)
39 Titi et al. [100] 2019 ✗ (0) High (3) Low (1) ✗ (0)
40 Vadillo et al. [101] 2013 Low (1) High (3) Medium (3) ✗ (0)
41 Villarreal et al. [102] 2014 Medium (2) Low (1) High (6) ✗ (0)
42 Xu et al. [105] 2017 ✗ (0) Low (1) Medium (3) ✗ (0)
43 Yu et al. [104] 2022 Medium (2) Low (1) High (5) ✗ (0)
44 Yu et al. [131] 2017 Low (1) High (3) Low (1) ✗ (0)
45 Zafeiropoulos et al. [107] 2024 Medium (2) High (3) Medium (3) High (2)
46 Zeshan et al. [132] 2023 Low (1) Medium (2) Medium (3) ✗ (0)
47 Zhang et al. [133] 2014 Low (1) Medium (2) Low (2) ✗ (0)
48 Zhou et al. [103] 2022 ✗ (0) ✗ (0) Low (1) ✗ (0)